OpenStudy (anonymous):

Explain how you would find the exact value of 3 over the square root of 8..

OpenStudy (mayankdevnani):

Can you write this problem in numeric way? @supanick14

OpenStudy (anonymous):

how

OpenStudy (mayankdevnani):

\[\huge \bf \frac{3}{\sqrt{8}}\] Is this your problem?

OpenStudy (mayankdevnani):

I am sure this is your problem

OpenStudy (dls):

\[\Huge \sqrt[8]{3}\] this is the problem?

OpenStudy (mayankdevnani):

now, we can write square root of 8 as :- \[\huge \bf \sqrt{8}=2\sqrt{2}\] now plug this value in your question,we get \[\huge \bf \frac{3}{2\sqrt{2}}\] and \[\huge \bf \sqrt{2}=1.41\] it is an approximate value Then, \[\huge \bf \frac{3}{2 \times 1.41}=\frac{3}{2.82}=1.063\] so, \[\huge \bf \color{red}{1.063}~is~your~answer.\]

OpenStudy (mayankdevnani):

understood? @supanick14

OpenStudy (mayankdevnani):

are you sure @DLS this is the problem which is given by you?

OpenStudy (mayankdevnani):

@DLS and @supanick14 proof of your question :- http://www.wolframalpha.com/input/?i=3+over+the+square+root+of+8..

OpenStudy (anonymous):

 = 3 ⁄ √8  = 3 ⁄ √(4 • 2)  = 3 ⁄ [(√4) • (√2) ]  = 3 ⁄ (2  • √2) ... this is OK  = 1.5 ⁄ √2 ... or this  = 3  • (√2) ⁄ (2  • √2  • √2) ... multiplying by: (√2) ⁄ √2   = 3  • (√2) ⁄ 4 ... with the denominator rationalized. Note:  ... √2 : is an irrational number since it can't be expressed as either a finite  ...   decimal number OR a ratio of integer values. The only way to determine the value of √2 accurately to a finite number of decimal places is to use infinite series of √2. The more series terms evaluated, the more accurate the result.

OpenStudy (mayankdevnani):

@supanick14 your question said that find the exact value.

OpenStudy (mayankdevnani):

understood my way? @supanick14 reply me fast

OpenStudy (anonymous):

ya

OpenStudy (mayankdevnani):

good

OpenStudy (cggurumanjunath):

1.063